OpenStudy (anonymous):

How to find asymtope for this x^2/(x-2)

jimthompson5910 (jim_thompson5910):

vertical: set the denominator equal to zero and solve for x oblique: use polynomial long division

OpenStudy (anonymous):

for vertical x=2 then for the second..

jimthompson5910 (jim_thompson5910):

good, the VA is x = 2

OpenStudy (anonymous):

If the degree (the largest exponent) of the denominator is bigger than the degree of the numerator, the horizontal asymptote is the x-axis (y = 0). If the degree of the numerator is bigger than the denominator, there is no horizontal asymptote. If the degrees of the numerator and denominator are the same, the horizontal asymptote equals the leading coefficient (the coefficient of the largest exponent) of the numerator divided by the leading coefficient of the denominator

OpenStudy (anonymous):

how i,m going to find the second asymptote x^2/(x-2)=x+2+( 4/x-2)

jimthompson5910 (jim_thompson5910):

the quotient is x+2 so the oblique asymptote is y = x+2

OpenStudy (anonymous):

how to get the oblique asymtope.. if there have other method beside using limit???

jimthompson5910 (jim_thompson5910):

use polynomial long division